40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1579     Entry time: Sun Sep 13 10:40:29 2015
Author: Antonio 
Type: DailyProgress 
Category: BEAT 
Subject: PLL measured noise 

Summary

========

The goal of the TCN experiment is to measure the TC noise. This requires to lower down the noise level that we have at

the PLL output. I have decided to take noise measurement today in order to have a reference from the level we are starting with.

We should start to implement changes in order to lower the noise at the output and keep monitoring it.

----------------------------------------------------------------------------------------------------------------------------------------------------------------------------

 

Description

=========

I report a plot of the PLL noise in Vrms/sqrt(Hz) as it is not clear yet how to convert the units in Hz/sqrt(Hz). From older Elogs ID 889,

I see that there is available a calibration for the IFR 2023a, and this depends on the input range. I am not sure on what the input range is.

The measurement has been taken in dirrent frequency range in order to have a better resolution. For now, when we are at this level

of noise it is not worth it.

 

Note:

I have tried to lock the PLL with a FM deviation which is less than 300kHz but it was not possible. There is too much noise.

 

Figure 1:

Target noise and noise taken sometimes in the past by Evan’s (in Hz/sqrt(Hz)).

 

Figure 2:

Current measurements in Vrms/sqrt(Hz);

Beat freq = ~ 65 MHz;

FM deviation = 400kHz

gain = 20;

note: something happened after the first measurement, but at this point is not so important.

 

Data:

Data are on the lab control/home/data/20150912_PLL_noise

 

 

/

Attachment 1: NoiseEvan.pdf  105 kB  | Hide | Hide all
NoiseEvan.pdf
Attachment 2: Noise_20150912_Vrms.pdf  47 kB  | Hide | Hide all
Noise_20150912_Vrms.pdf
ELOG V3.1.3-